AQuA-RAT train #66727 · Probability · Counting & chance
Three cannons are firing at a target
Three cannons are firing at a target. If their individual probabilities to hit the target are 0.2, 0.1, and 0.7 respectively, what is the probability that none of the cannons will hit the target after one round of fire? prob. that all the cannons will hit the target = .06 prob. that none of the cannons will hit = 1 - .06 = .94
AQuA-RAT — Algebra Question Answering with Rationales (Ling et al., DeepMind, 2017) · train #66727
Stuck? A gentle nudge
No hints for this one.
Your answer
Checked against the answer in AQuA-RAT.